OpenStudy (anonymous):

Belinda wants to invest $1000. The table below shows the value of her investment under two different options for three different years: Number of years 1 2 3 Option 1 (amount in dollars) 1300 1690 2197 Option 2 (amount in dollars) 1300 1600 1900 Part A: What type of function, linear or exponential, can be used to describe the value of the investment after a fixed number of years using option 1 and option 2? Explain your answer. (2 points) Part B: Write one function for each option to describe the value of the investment f(n), in dollars, after n years. (4 points) Part C: Beli

OpenStudy (anonymous):

Part C: Belinda wants to invest in an option that would help to increase her investment value by the greatest amount in 20 years. Will there be any significant difference in the value of Belinda's investment after 20 years if she uses option 2 over option 1? Explain your answer, and show the investment value after 20 years for each option. (4 points)

OpenStudy (anonymous):

I ONLY NEED PART C IGIVE MEDALS :)

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

@Kainui

OpenStudy (anonymous):

help please?

OpenStudy (danjs):

20 year value, n = 20, put that in both of your functions from part B

OpenStudy (anonymous):

can you show me how to solve it

OpenStudy (danjs):

what are your functions from part b ?

OpenStudy (anonymous):

1300+300n 1300+300n* 1.3n

OpenStudy (danjs):

hmm The first option is exponential form (1000)*(r)^n The second option is linear form 1000 + n*d have to figure what r and d are

OpenStudy (danjs):

Option 1) THe function looks like (initial investment)*(r)^n \[f(n) = 1000*(r)^n\] r is the common ratio of consecutive terms 1300/1000 = 1.3 or 1690/1300 = 1.3 or 2197/1690 = 1.3 r is 1.3

OpenStudy (danjs):

f(n) = 1000(1.3)^n

OpenStudy (danjs):

Each next term is the previous multiplied by 1.3

OpenStudy (danjs):

Option 2) Each term increases by the same amount over the previous term, it is always 300 more

OpenStudy (danjs):

f(n) = (starting value) + n*(common difference) f(n) = 1000 + 300*n

OpenStudy (anonymous):

so i use this to solve for the difference right?

OpenStudy (danjs):

1) f(n) = 1000(1.3)^n 2) f(n) = 1000 + 300*n yep, put n=20 in both options and calculate the values

OpenStudy (anonymous):

do i subtract both of the answers i get for the answer

OpenStudy (danjs):

you dont have to, just compare them

OpenStudy (danjs):

option 1 should be way larger

OpenStudy (anonymous):

yes it was that you so much :)

OpenStudy (anonymous):

*thank

OpenStudy (danjs):

welcome